# Multimodal Fusion on Low-quality Data: A Comprehensive Survey

### v2

@@ 3.2.3 Kernel learning based
Kernel learning based methods transform the learning phase from the raw-feature space into the kernel space. Similar to the graph learning based methods, each element of the kernel matrix denotes a kind of the similarity or connection values of the corresponding two samples. For incomplete multimodal data, the constructed kernels w.r.t. different modalities will also be incomplete with missing rows and columns. Thus, the target of kernel learning based methods for incomplete multimodal data is to explore the information of these incomplete kernels. In this branch, most methods consider kernel imputation as we mentioned above like model (3) while relatively few works focus on exploiting certain available kernel information. Late fusion incomplete multimodal clustering [100] is one of the representative kernel imputation-free methods for incomplete multimodal data. The method does not need to impute the missing kernel elements but tries to simultaneously impute the missing representations corresponding to the missing modalities and learn the consensus representation in a joint framework. Furthermore, to reduce the low-quality imputation of the latent representations, Liu et al[101] proposed an efficient and effective regularized incomplete multimodal clustering method (EERIMVC), which integrates the initialized consensus representation as prior knowledge. In this model, such prior representation is obtained by the conventional complete kernel algorithm with zero or average filling. Therefore, the quality of the initialized consensus representation will be an important factor to the performance. To break the gap between representation learning and label prediction, a one-stage incomplete multimodal clustering method is proposed [102]. Compared with EERIMVC, this method replaces the consensus representation with a clustering label matrix and an orthogonal matrix.
To sum up, for the incomplete multimodal raw data $\left\{{{X^{\left(m\right)}}\in{R^{{d_{m}}\times n}}}\right\}_{m=1}^{l}$ or derived incomplete kernel data $\left\{{{K^{\left(m\right)}}\in{R^{n\times n}}}\right\}_{m=1}^{l}$ with a binary matrix $W\in{\left\{{0,1}\right\}^{n\times l}}$ indicating the location of missing modalities, the late fusion based kernel methods generally seek to learn the consensus representation $P\in{R^{n\times c}}$ from the initialized representations $\left\{{{{\bar{P}}^{\left(m\right)}}\in{R^{{n_{m}}\times c}}}\right\}_{m=1}^{l}$ of the available modalities as follows:
$\begin{array}[]{l}\mathop{\max}\limits_{\phi\left({P,\left\{{{{\tilde{P}}^{\left(m\right)}},{Q^{\left(m\right)}}}\right\}_{m=1}^{l},\gamma}\right)}f\left({P,\sum\limits_{m=1}^{l}{\gamma_{m}^{2}{{\tilde{P}}^{\left(m\right)}}{Q^{\left(m\right)}}}}\right)\\ {\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}+\varphi\left({P,\left\{{{{\tilde{P}}^{\left(m\right)}},{Q^{\left(m\right)}}}\right\}_{m=1}^{l},\gamma}\right),\end{array}$ (7)
where function $f$ is generally set as a trace operation in the above-mentioned works. ${Q^{\left(m\right)}}$ can be regarded as the perturbation matrix of the $m$-th modality. $\phi\left({P,\left\{{{{\tilde{P}}^{\left(m\right)}},{Q^{\left(m\right)}}}\right\}_{m=1}^{l},\gamma}\right)$ denotes the boundary constraint for these variables while $\varphi\left({P,\left\{{{{\tilde{P}}^{\left(m\right)}},{Q^{\left(m\right)}}}\right\}_{m=1}^{l},\gamma}\right)$ are the regularization constraints for these variables. In many works, ${\tilde{P}^{\left(m\right)}}$ is the recovered complete representation for the $m$-th modality, where some rows of ${\tilde{P}^{\left(m\right)}}\in{R^{n\times c}}$ corresponding to the available modalities are fixed as the original initialized ${\bar{P}^{\left(m\right)}}$.
@@ 3.2.4 Deep learning based
For supervised learning tasks, a cross-partial multimodal network is proposed in [109], which can flexibly learn a common representation for arbitrary missing data and preserve both common and complementary information. For the more difficult task with incomplete labels, a deep double incomplete multimodal multi-label learning network [110] is proposed. This deep network mainly introduces two mask constraints w.r.t. the missing view locations and missing label locations, to reduce the negative impact of the missing labels and missing views. A good property of the method is that it can handle all kinds of incomplete data classification tasks including supervised/semi-supervised cases, complete/incomplete data, and complete/incomplete label cases. Some representative incomplete multimodal learning methods are summarized in Table 4.
@@ 3.3 Discussion
To sum up, the ways to handle incomplete data can be commonly divided into two categories, imputation-based incomplete multimodal learning and imputation-free incomplete multimodal learning. Methods that focus on imputing incomplete modalities can be classified into two types: one indirectly recovers the missing modalities through graph or kernel completion, and the other directly fills in the original data. For imputation-free incomplete multimodal learning methods, their design philosophy focuses more on utilizing available modalities to explore information and understand multimodal data. However, both types of methods for handling incomplete multimodal data face deeper challenges. For example, the quality assessment of imputed instances for missing modalities is commonly overlooked. Besides, using prior missing information to mask unknown modalities is itself difficult to bridge the information gap and information imbalance issues caused by modal missing.
TABLE III: A summary of some representative incomplete multimodal learning works.
@@
Representation learning Classification [109]
@@ 4.1.3 Architecture-based method
TABLE IV: A summary of representative balanced multimodal learning studies.
@@ 4.2 Quality-discrepancy methods
In addition to the issue of discrepancy in unimodal learning property, even though all modalities convey the same concepts, they contain varying amounts of information related to the target event or objects, resulting a discrepancy in unimodal data quality. Given the inherently greedy nature of deep neural networks, recent studies have further uncovered that multimodal models tend to rely heavily on a single modality with high quality that provides sufficient target-related information, while neglecting the others, leading to their under-optimized [8, 10]. As shown in Fig. 4, Peng et al. [8] empirically verified that the quality of unimodal representations in the multimodal model is worse than that in the corresponding unimodal models, despite the overall superior performance of the multimodal models. Moreover, this under-optimization of each modality is imbalanced, and the modality with high quality is less under-optimized. They claim that this imbalanced problem is caused by the model preference for the high-quality modality with better performance (i.e., the dominant modality). For instance, the audio modality in the curated sound-oriented dataset, VGGSound [130], is less under-optimized compared to the vision modality. Theoretical analysis conducted by Huang et al.  [10] has further confirmed that for the jointly trained multimodal late-fusion framework, the encoder networks can only learn a subset of modalities, and other modalities may fail to be discovered well. This finding is consistent with empirical observations, indicating that the imbalanced preference for different modalities issue needs to be addressed in multimodal learning. To tackle this problem brought by discrepancies in unimodal data quality, a series of studies have been proposed, encompassing learning-objective-based method, optimization-based method, architecture-based method and data-augmentation-based method.
@@ 4.2.2 Optimization-based method
Optimization-based methods focus on the back-propagation stage. These methods respectively pay attention to the magnitude and direction of unimodal gradient, to improve the learning of modality with lower quality. To slow down the learning of preferred modality with high quality and accordingly offer more training efforts to other modalities, Peng et al. [8] proposed an on-the-fly gradient modulation strategy, which dynamically monitors the contribution difference of various modalities to the final prediction during training and mitigates the gradient magnitude of dominant modality to focus more on other modalities. Inspired by this gradient magnitude modulation idea, Sun et al. [123] extended it into the graph-based models of rumor detection task, adaptively adjusting the magnitude of gradients to alleviate the preference for high-quality modality. Similarly, Fu et al. [124] extended the idea of gradient magnitude modulation into audio-visual video parsing task. They further proposed a modality-separated decision unit to well evaluate the unimodal prediction in the more complex audio-visual video parsing framework with more cross-modal interaction. Besides the gradient magnitude, the direction of unimodal gradient is also explored. Fan et al. [122] suggested that the slow-learning modality’s updating direction is severely disturbed by the preferred high-quality modality, making it hard to its training. Subsequently, they utilize prototypes, which are the centroids of each class within the representation space, to refine the updating direction, aiming to improve unimodal performance.
@@ 5 Dynamic Multimodal Fusion
Current multimodal fusion methods often depends on the assumption that multimodal data are of static quality, which does not always hold in real-world scenarios. Handling multimodal data with dynamically varying quality is an inevitable problem for multimodal intelligence systems. Due to unexpected environmental factors and sensor issues, some modalities may suffer from poor reliability and lack of task-specific information. Besides, the quality of different modalities are dynamically varying from scenarios as shown in Fig. 5. This phenomenon motivates a new multimodal learning paradigm, i.e., dynamical multimodal fusion, which aims to adapt to the dynamically changing quality of multimodal data and selectively integrate task-specific information. In this section, we focus on the challenges of dynamical multimodal fusion and categorize current advances in this literature into three main lines including heuristic, attention-based and uncertainty-aware dynamic fusion.
TABLE V: A summary of our taxonomy of dynamic multimodal fusion .
@@
Uncertainty-aware Entropy Mutual information Semantic segmentation [144]
@@ 5.1 Heuristic dynamic fusion
Toward achieving robust multimodal fusion, previous works heuristically devise dynamic fusion strategies. These dynamic fusion methods are derived from human experience and knowledge (e.g., thermal modality is more reliable than RGB in nighttime) about the task at hand and application scenarios.
As a concrete example, illumination conditions can serve as a criterion to introduce dynamic fusion. In around-the-clock applications, previous work identifies that the quality of RGB and thermal modalities are temporally varying in multispectral pedestrian detection task [133]. While the RGB modality tends to contain more useful information compared to the thermal modality under normal illumination conditions, this relationship can be reversed under low light or nighttime conditions where the thermal modality becomes more reliable than RGB. Based on this observation, the authors proposed to utilize an illumination aware fusion module that fuse features from the two modalities adaptively. Similarly, to enable adaptation to varying illumination conditions, Guan et al[132] proposed an illumination-aware weighting mechanism that evaluates the lighting of each input scene. These illumination weights are utilized to dynamically integrate the outputs of sub-networks specialized for day and night detection and segmentation. The proposed method can further be applied to pedestrian detection in security surveillance and autonomous driving systems.
@@
Some relevant works also utilize task-specific modules to achieve dynamic fusion. To highlight a few, Panda et al[135] introduced AdaMML, an adaptive multimodal learning in efficient video recognition. By dynamically selecting optimal input modalities per video segment, this work contributes to data-dependent multimodal selection for efficient video recognition. In this work, a lightweight policy network is devised to dynamically select optimal input modalities for each video segment. AdaMML utilizes Gumbel-Softmax sampling for differentiable training of discrete decisions. Similarly, Xue et al[137] proposed DynMM, a novel approach for dynamic multimodal fusion in deep learning. It introduces modality-level and fusion-level decisions, allowing adaptive selection of input modalities and fusion operations based on data characteristics.
@@ 5.2 Attention-based dynamic fusion
As one of the most advanced variants of self-attention, transformers (based on multi-head self-attention) are frequently used in recent multimodal fusion methods to enhance the expressive power of the model. Nagrani et al.  proposed Fusion Bottlenecks, a new transformer architecture that restricts cross-modal interactions to bottleneck tokens and mid-late fusion layers only [141]. Fusion Bottlenecks force the model to share only necessary information between modalities, avoiding expensive full pairwise attention. Key to the approach is the insight that full pairwise attention between all modalities’ tokens is often redundant, and condensing interactions through fusion bottlenecks provides an efficient alternative. Girdhar et al. leveraged a shared Transformer architecture with shared parameters to handle diverse vision modalities including images, videos, and single-view 3D data. The key idea is to fuse information from different modalities into a shared embedding representation by employing a self-attention mechanism to model cross-modality relationship both temporally and spatially [142]. Wang et al. identified that existing Transformer variants might dilute unimodal modal attention weights on multimodal tasks, leading to sub-optimal overall performance. To overcome this, they propose to dynamically identifies uninformative tokens and replace them with aggregated features from other modalities for effective fusion [143]. Unlike prior works that mainly focus on enhancing reliability of multimodal fusion during training time, Yang et al[yang2023test] proposed to utilizing attention to fully handle corrupted modalities in test time. A confidence-aware loss is devised to enhance credible predictions and mitigate potential influence of noise. The proposed method provides new insights into leveraging attention mechanisms for effective test-time adaption in multimodal fusion.
@@ 5.3 Uncertainty-aware dynamic fusion
where =$\mathcal{M}=\{\{b_{k}\}_{k=1}^{K},u\}$ is the multimodal brief mass, $\mathcal{M}^{1}=\{\{b_{k}^{1}\}_{k=1}^{K},u^{1}\},\mathcal{M}^{2}=\{\{b_{k}^{2}\}_{k=1}^{K},u^{2}\}$ are unimodal brief masses and $b^{1}_{k},b^{2}_{k}$ is the belief mass assigned for $k$-th class according to each modality. $u^{1},u^{2}$ and $u$ are unimodal and joint uncertainty respectively. Thanks to the property of D-S theory, the final prediction of TMC will depend on the modality with low uncertainty. And thus the classification robustness is improved by integrating evidence from each modality.
Inspired by TMC, researchers deployed SL-based uncertainty-aware multimodal fusion in many scenarios, including medical image classification [149], object detection [148, 157], semantic segmentation [2]. In whole-slide medical image classification task, by integrating evidences from modalities of multiple scales, they significantly improve the medical classification performance in autoimmune diseases detection and Fibroma classification tasks. Similarly, Chen et al[158] also leverage Dirichlet distribution to model the uncertainty of multimodal representations and then fuse the segmentation and classification representations adaptively.
Although TMC exhibits a great performance, Liu et al[150] identified its limitations, including inability to guarantee an overall decrease in uncertainty during the fusion of diverse views and neglecting of consistency across views. To address these issues, they proposed an improved opinion aggregation framework, ensuring consistency across views and achieving more reliable predictions. Xu et al[159] pointed out that an important characteristic of TMC is ”After integrating another opinion into the original opinion, the obtained uncertainty mass will be reduced”. They argued this since when incorporating an unreliable or conflicting opinion, the uncertainty should increase. To solve this, they proposed a conflictive opinion aggregation strategy and theoretically proved the uncertainty would increase for conflictive instances.
Entropy is the most natural and straightforward way to capture modal-wise uncertainty. Tian et al[144] proposed an entropy-based multimodal fusion method, named Uncertainty-aware Noisy-Or Multimodal Fusion (UNO). Specifically, UNO first introduces several uncertainty measurements, including predictive entropy, mutual information based on Monte Carlo Dropout, as well as deterministic entropy. It then proceeds to combine these uncertainty metrics for fusion by selecting the most conservative (highest uncertainty) measurement and weighting the outputs from each modalities based on the uncertainty. Finally, a simple but effective noisy-or fusion scheme is deployed to fusing decisions from multiple modalities. In multimodal semantic segmentation task, UNO can be used to improve the robustness in the presence of various unknown input degradation.
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation operator, $\mathbf{h}^{1},\cdots,\mathbf{h}^{M}$ and $\mathbf{h}$ are unimodal and multimodal representations respectively. $conf^{m}$ is the predicted confidence for the $m$-th modality which measures the informativeness. Zhang et al[33] proposed a confidence-aware fusion method to selectively integrate features from RGB and thermal modalities for multi-spectral pedestrian detection. It predicts the classification confidence of each modality, then performs feature re-weighting to emphasize cues from more reliable modalityies as well as suppress the less useful ones based on the confidence scores.

### v3

@@ 3.2.3 Kernel learning based
Kernel learning based methods transform the learning phase from the raw-feature space into the kernel space. Similar to the graph learning based methods, each element of the kernel matrix denotes a kind of the similarity or connection values of the corresponding two samples. For incomplete multimodal data, the constructed kernels w.r.t. different modalities will also be incomplete with missing rows and columns. Thus, the target of kernel learning based methods for incomplete multimodal data is to explore the information of these incomplete kernels. Late fusion incomplete multimodal clustering [91] is one of the representative kernel imputation-free methods for incomplete multimodal data. The method does not need to impute the missing kernel elements but tries to simultaneously impute the missing representations corresponding to the missing modalities and learn the consensus representation in a joint framework. Furthermore, to reduce the low-quality imputation of the latent representations, Liu et al[92] proposed an efficient and effective regularized incomplete multimodal clustering method (EERIMVC), which integrates the initialized consensus representation as prior knowledge. In this work, such prior representation is obtained by the conventional complete kernel algorithm with zero or average filling. Therefore, the quality of the initialized consensus representation will be an important factor to the performance. To break the gap between representation learning and label prediction, a one-stage incomplete multimodal clustering method is proposed [93]. Compared with EERIMVC, this method replaces the consensus representation with a clustering label matrix and an orthogonal matrix.
@@ 3.2.4 Deep learning based
For supervised learning tasks, a incomplete multimodal learning framework is proposed in [100], which can flexibly learn a common representation for arbitrary missing data and preserve both common and complementary information. For the more difficult task with incomplete labels, a deep double incomplete multimodal multi-label learning network [101] is proposed. This deep network mainly introduces two mask constraints w.r.t. the missing view locations and missing label locations, to reduce the negative impact of the missing labels and missing views. A good property of the method is that it can handle all kinds of incomplete data classification tasks including supervised/semi-supervised cases, complete/incomplete data, and complete/incomplete label cases. More recent work [102] focuses on an even more generalized scenario than incomplete multimodal learning which is so called train-test modality mismatch, where the test multimodal data can contains more or less modalities than training data. Specifically, all modalities are available at pretraining stage, while a completely different subset is encountered by the model during supervised training and testing in downstream tasks.
In medical image processing, Sanaz et al. proposed MMCFormer [103], a novel missing modality compensation network to address the incomplete modalities for brain tumor segmentation. The proposed method is build upon 3D efficient transformer blocks and uses a co-training strategy to effectively train a missing modality network by ensuring feature consistency between complete modalities and incomplete modalities. Yao et al. proposed DrFuse [104] for effective medical diagnosis and forecasting prognosis with incomplete multimodal data. DrFuse first learns the common features shared between all modalities and modality-specific features within each separated modality and then weight and then combining the prediction from each modality to make the final prediction. Specifically, a disease-wise attention layer is devised to produce the patient- and disease-wise weights to tackle the modality inconsistency issue.
Some representative incomplete multimodal learning methods are summarized in Table 4.
@@ 3.3 Discussion
To sum up, the ways to handle incomplete data can be commonly divided into two categories, imputation-based incomplete multimodal learning and imputation-free incomplete multimodal learning. Methods that focus on imputing incomplete modalities can be classified into two types: one indirectly recovers the missing modalities through graph or kernel completion, and the other directly fills in the original data. For imputation-free incomplete multimodal learning methods, their design philosophy focuses more on utilizing available modalities to explore information and understand multimodal data. However, both types of methods for handling incomplete multimodal data face deeper challenges. For example, the quality assessment of imputed instances for missing modalities is commonly overlooked. Besides, using prior missing information to mask unknown modalities is itself difficult to bridge the information gap and information imbalance issues caused by modal missing. Moreover, incomplete cases such as unpaired views [105, 106, 107] and partially missing labels [108] also need more attention.
TABLE II: A summary of some representative incomplete multimodal learning works.
@@
Representation learning strategy Classification [100, 102, 104]
@@
Representation learning strategy Medical image segmentation [103]
@@ 4.1.3 Architecture-based method
TABLE III: A summary of representative balanced multimodal learning studies.
@@ 4.2 Quality-discrepancy methods
In addition to the issue of discrepancy in unimodal learning property, even though all modalities convey the same concepts, they contain varying amounts of information related to the target event or objects, resulting a discrepancy in unimodal data quality. Given the inherently greedy nature of deep neural networks, recent studies have further uncovered that multimodal models tend to rely heavily on a single modality with high quality that provides sufficient target-related information, while neglecting the others, leading to their under-optimized [8, 10]. As shown in Fig. 4, Peng et al. [8, 128] empirically verified that the quality of unimodal representations in the multimodal model is worse than that in the corresponding unimodal models, despite the overall superior performance of the multimodal models. Moreover, this under-optimization of each modality is imbalanced, and the modality with high quality is less under-optimized. They claim that this imbalanced problem is caused by the model preference for the high-quality modality with better performance (i.e., the dominant modality). For instance, the audio modality in the curated sound-oriented dataset, VGGSound [129], is less under-optimized compared to the vision modality. Theoretical analysis conducted by Huang et al.  [10] has further confirmed that for the jointly trained multimodal late-fusion framework, the encoder networks can only learn a subset of modalities, and other modalities may fail to be discovered well. This finding is consistent with empirical observations, indicating that the imbalanced preference for different modalities issue needs to be addressed in multimodal learning. To tackle this problem brought by discrepancies in unimodal data quality, a series of studies have been proposed, encompassing learning-objective-based method, optimization-based method, architecture-based method and data-augmentation-based method.
@@ 4.2.2 Optimization-based method
Optimization-based methods focus on the back-propagation stage. These methods respectively pay attention to the magnitude and direction of unimodal gradient, to improve the learning of modality with lower quality. To slow down the learning of preferred modality with high quality and accordingly offer more training efforts to other modalities, Peng and Wei et al. [8, 128] proposed an on-the-fly gradient modulation strategy, which dynamically monitors the contribution difference of various modalities to the final prediction during training and mitigates the gradient magnitude of dominant modality to focus more on other modalities. Inspired by this gradient magnitude modulation idea, Sun et al. [121] extended it into the graph-based models of rumor detection task, adaptively adjusting the magnitude of gradients to alleviate the preference for high-quality modality. Similarly, Fu et al. [122] extended the idea of gradient magnitude modulation into audio-visual video parsing task. They further proposed a modality-separated decision unit to well evaluate the unimodal prediction in the more complex audio-visual video parsing framework with more cross-modal interaction. Besides the gradient magnitude, the direction of unimodal gradient is also explored. Fan et al. [120] suggested that the slow-learning modality’s updating direction is severely disturbed by the preferred high-quality modality, making it hard to its training. Subsequently, they utilize prototypes, which are the centroids of each class within the representation space, to refine the updating direction, aiming to improve unimodal performance.
@@ 5 Dynamic Multimodal Fusion
Due to unexpected environmental factors and sensor issues, some modalities may suffer from poor reliability and lack of task-specific information. Besides, the quality of different modalities are dynamically varying from scenarios. We borrow a case from [131], as shown in Fig. 5, thermal modality tends to provide more task-specific information compared to RGB modality during nighttime, whereas the situation is reversed during the daytime. This phenomenon motivates a new multimodal learning paradigm, i.e., dynamical multimodal fusion, which aims to adapt to the dynamically changing quality of multimodal data and selectively integrate task-specific information. In this section, we focus on the challenges of dynamical multimodal fusion and categorize current advances in this literature into three main lines including heuristic, attention-based and uncertainty-aware dynamic fusion.
TABLE IV: A summary of our taxonomy of dynamic multimodal fusion .
@@
Uncertainty-aware Entropy Classification [143]
Semantic segmentation [144]
@@ 5.1 Heuristic dynamic fusion
Toward achieving robust multimodal fusion, previous works heuristically devise dynamic fusion strategies. These dynamic fusion methods are derived from human experience and knowledge (e.g., thermal modality is more reliable than RGB in nighttime) about the task at hand and application scenarios. As a concrete example, illumination conditions can serve as a criterion to introduce dynamic fusion. In around-the-clock applications, previous work identifies that the quality of RGB and thermal modalities are temporally varying in multispectral pedestrian detection task [131]. While the RGB modality tends to contain more useful information compared to the thermal modality under normal illumination conditions, this relationship can be reversed under low light or nighttime conditions where the thermal modality becomes more reliable than RGB. Based on this observation, the authors proposed to utilize an illumination aware fusion module that fuse features from the two modalities adaptively. Similarly, to enable adaptation to varying illumination conditions, Guan et al[132] proposed an illumination-aware weighting mechanism that evaluates the lighting of each input scene. These illumination weights are utilized to dynamically integrate the outputs of sub-networks specialized for day and night detection and segmentation. The proposed method can further be applied to pedestrian detection in security surveillance and autonomous driving systems.
@@
Some relevant works also utilize task-specific modules to achieve dynamic fusion. To highlight a few, Panda et al[134] introduced AdaMML, an adaptive multimodal learning in efficient video recognition. By dynamically selecting optimal input modalities per video segment, this work contributes to data-dependent multimodal selection for efficient video recognition. In this work, a lightweight policy network is devised to dynamically select optimal input modalities for each video segment. AdaMML utilizes Gumbel-Softmax sampling for differentiable training of discrete decisions. Similarly, Xue et al[136] proposed DynMM, a novel approach for dynamic multimodal fusion. It introduces modality-level and fusion-level decisions, allowing adaptive selection of input modalities and fusion operations based on data characteristics.
@@ 5.2 Attention-based dynamic fusion
As one of the most advanced variants of self-attention, transformers (based on multi-head self-attention) are frequently used in recent multimodal fusion methods to enhance the expressive power of the model. Nagrani et al.  proposed Fusion Bottlenecks, a new transformer architecture that restricts cross-modal interactions to bottleneck tokens and mid-late fusion layers only [140]. Fusion Bottlenecks force the model to share only necessary information between modalities, avoiding expensive full pairwise attention. Key to the approach is the insight that full pairwise attention between all modalities’ tokens is often redundant, and condensing interactions through fusion bottlenecks provides an efficient alternative. Girdhar et al. leveraged a shared Transformer architecture with shared parameters to handle diverse vision modalities including images, videos, and single-view 3D data. The key idea is to fuse information from different modalities into a shared embedding representation by employing a self-attention mechanism to model cross-modality relationship both temporally and spatially [141]. Wang et al. identified that existing Transformer variants might dilute unimodal modal attention weights on multimodal tasks, leading to sub-optimal overall performance. To overcome this, they propose to dynamically identifies uninformative tokens and replace them with aggregated features from other modalities for effective fusion [142]. Unlike prior works that mainly focus on enhancing reliability of multimodal fusion during training time, Yang et al[157] proposed to utilizing attention to fully handle corrupted modalities in test time. A confidence-aware loss is devised to enhance credible predictions and mitigate potential influence of noise.
@@ 5.3 Uncertainty-aware dynamic fusion
where =$\mathcal{M}=\{\{b_{k}\}_{k=1}^{K},u\}$ is the multimodal brief mass, $\mathcal{M}^{1}=\{\{b_{k}^{1}\}_{k=1}^{K},u^{1}\},\mathcal{M}^{2}=\{\{b_{k}^{2}\}_{k=1}^{K},u^{2}\}$ are unimodal brief masses and $b^{1}_{k},b^{2}_{k}$ is the belief mass assigned for $k$-th class according to each modality. $u^{1},u^{2}$ and $u$ are unimodal and joint uncertainty respectively. Thanks to the property of D-S theory, the final prediction of TMC will depend on the modality with low uncertainty. And thus the classification robustness is improved by integrating evidence from each modality. Inspired by TMC, researchers deployed SL-based uncertainty-aware multimodal fusion in many scenarios, including medical image classification [149], object detection [148, 158], semantic segmentation [2]. In whole-slide medical image classification task, by integrating evidences from modalities of multiple scales, they significantly improve the medical classification performance in autoimmune diseases detection and Fibroma classification tasks. Similarly, Chen et al[159] also leverage Dirichlet distribution to model the uncertainty of multimodal representations and then fuse the segmentation and classification representations adaptively. Although TMC exhibits a great performance, Liu et al[150] identified its limitations, including inability to guarantee an overall decrease in uncertainty during the fusion of diverse views and neglecting of consistency across views. To address these issues, they proposed an improved opinion aggregation framework, ensuring consistency across views and achieving more reliable predictions.
Entropy is the most natural and straightforward way to capture modal-wise uncertainty. Tian et al[144] proposed an entropy-based multimodal fusion method, named Uncertainty-aware Noisy-Or Multimodal Fusion (UNO). Specifically, UNO first introduces several uncertainty measurements, including predictive entropy, mutual information based on Monte Carlo Dropout, as well as deterministic entropy. It then proceeds to combine these uncertainty metrics for fusion by selecting the most conservative (highest uncertainty) measurement and weighting the outputs from each modalities based on the uncertainty. Finally, a simple but effective noisy-or fusion scheme is deployed to fusing decisions from multiple modalities. In multimodal semantic segmentation task, UNO can be used to improve the robustness in the presence of various unknown input degradation. Similarly, Zhang et al. [143] proposed to leverage the entropy of prediction as a proxy to gauge the importance of each modality, and then dynamically weight and fuse the unimodal predictions. By introducing such test-time dynamic fusion mechanism, the proposed method can be aware of uncertainty raised upon the varying quality of different modalities, and further enhance the effectiveness of multimodal fusion.
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation operator, $\mathbf{h}^{1},\cdots,\mathbf{h}^{M}$ and $\mathbf{h}$ are unimodal and multimodal representations respectively. $conf^{m}$ is the predicted confidence for the $m$-th modality which measures the informativeness. Zhang et al[26] proposed a confidence-aware fusion method to selectively integrate features from RGB and thermal modalities for multi-spectral pedestrian detection. It predicts the classification confidence of each modality, then performs feature re-weighting to emphasize cues from more reliable modalities as well as suppress the less useful ones based on the confidence scores.